OpenStudy (anonymous):

Solve the equation on the interval [0, 2π). tan2x sin x = tan2x A. 0, (pi)/2 B. (pi)/2 , 2π C. (pi)/2 , π D. 0, π

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

hint: tan2x sin x = tan2x tan2x sin x - tan2x = 0 tan(2x) ( sin(x) - 1) = 0 tan(2x) = 0 or sin(x) - 1 = 0

OpenStudy (anonymous):

i dont understand step 3.. @jim_thompson5910

jimthompson5910 (jim_thompson5910):

I factored out tan(2x) using the distributive property

jimthompson5910 (jim_thompson5910):

Notice if you distribute tan(2x) back in, you'll get step 2 again

OpenStudy (anonymous):

ohh. ok! so 0 is one of the choices. how do I figure out which one is the second choice?

jimthompson5910 (jim_thompson5910):

Solve sin(x) - 1 = 0 for x and tell me what you get

OpenStudy (anonymous):

pi/2!

OpenStudy (anonymous):

so its A?

jimthompson5910 (jim_thompson5910):

you got it

OpenStudy (anonymous):

thanks!

jimthompson5910 (jim_thompson5910):

np